A survey was conducted that asked 1011 people how many books they had read in the past year. Results indicated that

x overbarx equals=13.6

books and s equals=16.6

books.

Construct a 90​% confidence interval for the mean number of books people read. Interpret the interval.

Construct a 90​% confidence interval for the mean number of books people read and interpret the result.
Answer #1

Given that,

\bar x = 13.6

s =16.6

n = 1011

Degrees of freedom = df = n - 1 =1011 - 1 = 1010

At 90% confidence level the t is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.1

\alpha / 2 = 0.1 / 2 = 0.05

t\alpha /2,df = t0.05,1010 =1.646

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.646* (16.6 / \sqrt 1011) = 0.86

The 90% confidence interval is,

\bar x - E < \mu < \bar x + E

13.6 - 0.86< \mu < 13.6+ 0.86

12.74 < \mu < 14.46

(12.74,14.46 )

LOWER LIMIT 12.74 TO

UPPER LIMIT 14.46